OpenStudy (anonymous):

Find the curvature k of r(t) = <2e^(-2t), -2cos(10t), -8t> at the point (2, -2, 0)

OpenStudy (anonymous):

is this right for the first part?? \[(\sqrt{2,560,000\cos 10t^2 + 4096e^(-4t) + 25600e^(-4t) (5\cos(10t) + \sin(10t)^2 ) } / (\sqrt{16e^(-4t) + 40,000\sin(10t)^2 + 64})\]

OpenStudy (anonymous):

er wow jesus

OpenStudy (anonymous):

bleh... ok so never mind you shouldn't have to do all of that.. it should be evaluated at r(0) right?